OpenStudy (anonymous):

which number equals sqrt1,377 51 3sqrt17 17sqrt3 9sqrt17

OpenStudy (anonymous):

\( \huge 1377=81*17 \) now u tell which is correct?

OpenStudy (anonymous):

\[17\sqrt{3}\]

OpenStudy (anonymous):

notice that \[\huge \sqrt{1377} = \sqrt{81*17}=\sqrt{81}*\sqrt{17}\]

OpenStudy (anonymous):

\[\huge \sqrt{81}=9\]

OpenStudy (anonymous):

sqrt9*9*17 9sqrt17